# Conceptualisation, Methodology, Software, Validation, Writing - original draft

### v2

@@ Introduction
The influence of network complexity, phase configuration, and per-node connection on classification accuracy was first analysed. The per-node connection is the number of households connected to a point, e.g., a multi-unit building. Given datasets with SM errors but without available recordings of user-feeder connections, Table 3 presents the accuracy comparison under five modified configurations of LVDNs: a network with two complex-structured feeders (CN1), a network with a simple-structured and one complex-structured feeder (CN2), a network with two simple-structured feeders (SN), network with longer simple-structured feeders (SNL), and a network with two simple-structured feeders with more branches (SNB). Another complex-structured feeder or simple-structured feeder is added to the above five LVDNs to construct LVDNs with three feeders (as shown in the fifth column). The main difference between a complex network and a simple network is the length of feeders and the number of branches. The topology of CN2 and SNB are shown in Fig. 8.
Figure 9: PCC in first row and FPCC in second row among voltage datasets: a) with single-phase customers and b) with three-phase customers
Table 4: Average accuracy under three-phase datasets and multiple connections at nodes with SM errors
@@
The PCC and MFP among customers in CN1 with multiple connections per node are depicted in Fig. 9. As illustrated in Fig. 9(a), the expression in (7) reduces the voltage correlation between users connected to different feeders. This step helps distinguish between different connections more clearly. However, the complexity of the voltage data correlation shown in Fig. 9(b)—reflected in the similarity of the correlation distribution — makes it extremely challenging to simultaneously identify both the phase label and the user-feeder connection using only voltage data. A balanced LVDN represents a scenario where all users are connected to the same phase (i.e., a single-phase LVDN), whereas an unbalanced LVDN corresponds to a three-phase configuration.
@@
Table 4 indicates that the traditional KNN algorithm consistently outperforms the proposed modified KNN approach when the parameter $k$ is less than 8. However, the modified KNN demonstrates improved robustness as the proportion of missing data increases, particularly at higher unknown rates and a larger $k$. Specifically, the modified KNN method outperforms traditional KNN at higher levels of unknown information (e.g., 50%), achieving accuracy levels exceeding 87%. This suggests the proposed method may better preserve identification accuracy under limited accurate recordings from DSOs. In the simple network with branches (SNB), the performance gap between the two methods becomes less pronounced. Traditional KNN shows a slight advantage at intermediate unknown rates (e.g., 15%–20%) and smaller $k$ values, whereas the modified KNN achieves comparable or superior performance at higher levels of missing data and larger $k$ values. Fig. 10 depicts the accuracy of the approaches under datasets with various missing data percentages. The accuracy declines slightly by about 5% as the percentage of missing data increases, even at 70% missing data. This demonstrates the strong robustness of the proposed method against missing data.
Figure 10: Accuracy under incomplete dataset: a) modified KNN with 20% unknown labels, b) traditional KNN with 20% unknown label, c) modified KNN with 50% unknown labels, d) traditional KNN with 50% unknown labels.
@@
The influence of feeder complexity, environmental conditions, and load distribution on identification accuracy is analyzed, with results summarized in Table 5 and 6. CC_PV: Represents a complex-structured feeder from CN1, with photovoltaics (PV) installed at each house. SC_PV: Corresponds to a simple-structured feeder from SN, also with PV at each house. CC and SC: Represent feeders from CN1 and SN with PV, but with datasets preprocessed using the method described in Section 2.4. These four feeders serve single-phase users only.CC_3_NU and SC_3_NU: Denote complex and simple feeders equipped with three-phase meters, where the load is non-uniformly distributed among the three phases. “+20” indicates an additional 20 customers connected to the feeder. “Long feeder” refers to an extended feeder length compared to the original configuration. Accuracy and clustering purity are used to evaluate the performance of phase identification, as formulated in the expression (15).
@@
Table 5: Average purity (%) under datasets form feeders with different components

### v3

@@ Introduction
The influence of network complexity, phase configuration, and per-node connection on classification accuracy was first analysed. The per-node connection is the number of households connected to a point, e.g., a multi-unit building. The test networks are used to represent LVDNs with varying structural characteristics in terms of feeder length and branching complexity. Five representative configurations are considered, as summarised in Table 3: CN1, CN2, SN, SNL, and SNB. The topology of CN2 and SNB is shown in Fig. 8.
In this work, feeder structures are classified as simple or complex based on two key topological attributes: feeder length and the number of branches. Specifically, simple-structured feeders are characterised by relatively short radial paths with limited branching, while complex-structured feeders exhibit longer feeder lengths and multiple branching points, resulting in deeper topologies. In Fig. 8. The red diamond markers represent connection nodes, i.e., the electrical connection points between the main feeder and end-users. The SNB represents the SN with a higher number of lateral connections originating from the main cable, i.e., more branch points where users are connected.
Given datasets with SM errors but without available recordings of user-feeder connections, Table 4 presents the accuracy comparison under five modified configurations of LVDNs.
Figure 9: PCC in first row and FPCC in second row among voltage datasets: a) with single-phase customers and b) with three-phase customers.
Table 5: Average accuracy under three-phase datasets and multiple connections at nodes with SM errors.
@@
The PCC and MFP among customers in CN1 with multiple connections per node are depicted in Fig. 9. As illustrated in Fig. 9(a), the expression in (7) reduces the voltage correlation between users connected to different feeders. This step helps distinguish between different connections more clearly. However, the complexity of the voltage data correlation shown in Fig. 9(b) makes it extremely challenging to simultaneously identify both the phase label and the user-feeder connection using only voltage data. A balanced LVDN represents a scenario where all users are connected to the same phase (i.e., a single-phase LVDN), whereas an unbalanced LVDN corresponds to a three-phase configuration.
@@
It is observed that the identification accuracy of the proposed method declines when applied to three-phase networks featuring three or more outgoing cables. This performance degradation highlights an inherent limitation of relying solely on voltage profiles in scenarios with high branch density and complex coupling. To overcome this shortcoming, incorporating multi-source datasets represents a promising avenue for future enhancement. Leveraging such auxiliary information can expand the feature space, thereby significantly improving the discriminative power and spatial resolution of the algorithm in more sophisticated grid structures.
@@
Table 5 shows that the traditional Euclidean distance-based KNN consistently outperforms the proposed modified KNN when the neighborhood size is small (e.g., $k<8$), particularly under low levels of missing data. This behavior is expected, as a smaller $k$ relies on highly local information, where the standard distance metric can effectively capture similarity when sufficient accurate observations are available. However, this advantage diminishes as the proportion of missing data increases and the neighborhood size becomes larger. Under such conditions, the modified KNN demonstrates improved accuracy, particularly at higher unknown rates (e.g., 50%) and larger $k$, where it achieves accuracy levels exceeding 87%. This indicates that the proposed distance formulation is more effective in mitigating the distortion introduced by incomplete measurements, thereby preserving identification performance when reliable neighbor information becomes sparse. It is important to note that the selection of $k$ is inherently data-dependent in practical DSO applications. In data-sparse scenarios—where only a limited number of neighboring users have reliable phase labels—a small $k$ (e.g., k=1k=133) is typically preferred, in which case the traditional KNN may be sufficient. In contrast, in more realistic settings with larger available datasets, DSOs often rely on a larger neighborhood to improve statistical reliability. In such cases, the proposed modified KNN provides clear advantages in terms of robustness and stability. This trend is further confirmed in the SNB network, where increased branching introduces additional structural complexity. While the traditional KNN shows a slight advantage at intermediate unknown rates (e.g., 15%–20%) and small $k$, the modified KNN achieves comparable or superior performance as missing data increases or when larger neighborhoods are considered. Overall, these results suggest that the proposed method is not intended to replace traditional KNN in all scenarios, but rather to enhance accuracy and options under realistic conditions characterized by uncertain neighborhood information. This flexibility allows DSOs to select or adapt the method based on data availability and operational requirements.
Figure 10: Mean accuracy and corresponding performance gap under varying levels of data incompleteness ($k=3$).
To validate the robustness of the proposed method with respect to network topology, missing data, and measurement noise, a comprehensive set of experiments was conducted under varying scenarios. Specifically, performance was evaluated across different levels of incompleteness of the data, varying proportions of available labeled data, and different noise magnitudes. The missing data ratio varied from $10\%$ to $90\%$. Regarding SM noise, four classes of SMs were considered according to IEC 62053-21, corresponding to error levels of $0.2\%$, $0.5\%$, $1\%$, and $2\%$ (extended to $5\%$ for extreme scenarios). Besides, to compare the accuracy between the traditional KNN and the improved KNN under different missing data conditions, an accuracy gap ($\Delta$ accuracy) was introduced, defined as the difference between the average accuracy of the modified KNN algorithm and that of the traditional KNN algorithm. In this analysis, the number of neighbours was fixed at $k=3$. The corresponding results are illustrated in Fig. 10.
From the line plots in Fig. 10, it can be observed that the accuracy of the proposed method remains relatively stable across different levels of missing data, indicating strong robustness against data incompleteness. Furthermore, the bar charts show that the accuracy gap is generally negative under the considered noise levels when $k=3$, suggesting that the traditional KNN slightly outperforms the improved method in these scenarios. Nevertheless, the performance difference remains within $10\%$, which is consistent with the observations reported in Table 4.
Fig. 11 presents the comparison under different noise conditions. As the noise level increases, the accuracy of the traditional KNN gradually decreases, indicating its sensitivity to measurement noise due to its reliance on Euclidean distance. In contrast, the proposed method exhibits a different trend. The accuracy initially slightly increases and then decreases slightly as noise grows, while maintaining comparable overall performance. This is because the proposed method is based on correlation analysis rather than Euclidean distance, making it inherently more robust to noise. Moderate magnitude noise can enhance the distinguishability between different users due to heterogeneous noise, leading to a temporary improvement in performance. However, as the noise magnitude becomes large, the similarity induced by high-amplitude Gaussian noise impacts correlation coefficient calculation, leading to accuracy reduction.
Based on Table 4 and Fig. 1011, it can be concluded that the two methods exhibit complementary characteristics under different conditions. For scenarios involving high noise levels or larger $k$ values, the improved KNN tends to achieve better performance. Conversely, under low-noise conditions and smaller $k$, the traditional KNN can serve as a competitive alternative. Therefore, an adaptive selection between the two methods can be employed to ensure reliable accuracy across diverse operating conditions.
Figure 11: Mean accuracy under multiple datasets with different noise level.
@@
The influence of feeder complexity, environmental conditions, and load distribution on identification accuracy is analyzed, with results summarized in Table 6 and 7. CC_PV: Represents a complex-structured feeder from CN1, with photovoltaics (PV) installed at each house. SC_PV: Corresponds to a simple-structured feeder from SN, also with PV at each house. CC and SC: Represent feeders from CN1 and SN with PV, but with datasets preprocessed using the method described in Section 2.4. These four feeders serve single-phase users only.CC_3_NU and SC_3_NU: Denote complex and simple feeders equipped with three-phase meters, where the load is non-uniformly distributed among the three phases. “+20” indicates an additional 20 customers connected to the feeder. “Long feeder” refers to an extended feeder length compared to the original configuration. Accuracy and clustering purity are used to evaluate the performance of phase identification, as formulated in the expression (15).
@@
Table 6: Average purity (%) under datasets form feeders with different components.
@@
Limitation analysis
Based on the above discussion and simulation result analysis, it is verified that the proposed method offers an effective solution for DSOs to maintain accurate topology within Digital Twin frameworks via non-intrusive voltage datasets. However, its broad application faces three potential challenges:
  • Network Topology Complexity: The identification performance is increasingly challenged by sophisticated distribution structures, particularly in networks featuring more than three cables (e.g., high branch density), which complicates the correlation patterns.
  • Data quality heterogeneity: according to the robustness analysis in Section 3.2, inconsistent SMs precision poses difficulties in tuning the hyperparameter $k$.
  • Data granularity: if users only share the maximum/minimum voltage value [liu2025topology], the application of the developed approach will be significantly limited.
  • Evolving load dynamics: The proliferation of heat pumps may change the distribution of winter data, potentially introducing new patterns of similarity that challenge the feasibility of the proposed approach.
In future work, these limitations will be further addressed. Meanwhile, the ongoing transition toward advanced smart meters (SMs) with enhanced sensing and communication capabilities provides significant opportunities to further improve the accuracy and practical applicability of the proposed approach in increasingly digitised power grids.